OpenStudy (anonymous):

The perimeter of a rectangle is forty-eight feet. If the length measures sixteen feet, which is the width? 5 feet 8 feet ten feet sixteen feet

OpenStudy (31356):

8

OpenStudy (31356):

16+16=32 48-32=16 16/2=8

OpenStudy (31356):

Medal? If helped....

OpenStudy (31356):

Thanks!